a 0.910 g sample of magnesium was reacted with oxygen to form 1.510 g of magnesium oxide.how much magnesium is needed to produce 0.850 g of magnesium oxide? how much magnesium and oxygen are needed to produce exactly 2.55 of magnesium oxide
1
Expert's answer
2021-10-11T06:12:29-0400

In the first case we can calculate the Mass of oxygen used =1.510g-0.910g


For the second question;

"2Mg+O_2\to 2MgO"

now moles of magnesium oxide "=\frac{0.850}{2(24+16)}=0.0106"

mole ratio; "2Mg:2MgO=2:2=1:1"

moles of Mg=0.0106

mass of magnesium"=48\times 0.0106=0.5088g"


For the next question;

"2Mg+O_2\to 2MgO"

mass of magnesium oxide "=\frac{2.55}{2(24+16)}=0.031875"

mole ratio "2MgO:2Mg=2:2=1:1"

mass of Mg "=2\times 24\times 0.031875=1.53g"

mass of oxygen=2.55g-1.53g=1.02g